# Impact of memory-burdened black holes on primordial gravitational waves in light of Pulsar Timing Array

### v1

@@ I Introduction
The theoretical and phenomenological framework outlined above relies on a semiclassical treatment of black-hole evaporation, which has been recently challenged [40, 41, 42]. Hawking’s semiclassical calculations for 4-dimensional black holes are expected to be invalid when the black hole has lost approximately half of its initial mass. This is primarily due to the neglection of a back-reaction, wherein the quantum information retained in the system’s memory (or an excess of entropy) exerts a reciprocal effect on the black hole. This effect, also known as the “memory burden”, leads to a significant extension of the black hole’s lifetime due to the accumulation of this quantum information or memory, within the black hole. Such a pivotal deviation of the PBH evaporation from the semiclassical approximation yields profound phenomenological consequences which have been investigated in the context of dark matter [43, 44, 45, 46, 47, 48, 49, 50] and primordial GWs sourced by PBHs [51, 52, 53, 54, 55, 56, 49, 57].
@@ II Impact of memory burden on PBH evaporation dynamics
where $\mathcal{A}=\mathcal{G}\,g_{*}(T_{\rm PBH})/({30720\pi})$ with $\mathcal{G}\simeq 3.8$ being the greybody factor, $g_{*}(T_{\rm PBH})\simeq 108$ is the number of relativistic degrees of freedom below the Hawking temperature $T_{\rm PBH}$, and $S(M_{\rm PBH})=4\pi M_{\rm PBH}^{2}/M_{\rm Pl}^{2}$ with $M_{\rm Pl}=1.22\times 10^{19}~{}{\rm GeV}$ is the black-hole entropy. When memory burden sets in at $t_{\rm mb}=(1-q^{3})t_{\rm ev}^{\rm sc}$, where $t_{\rm ev}^{\rm sc}=M_{\rm PBH}^{3}/(3\mathcal{A}\,M_{\rm Pl}^{4})$ is the semiclassical evaporation time, the evaporation dynamics starts to be modified allowing for a prolonged PBH lifetime before complete evaporation. The total evaporation time $t_{\rm ev}$ can then be split up into,
@@
where $\tilde{t}_{\rm ev}$ is the time from when memory burden effects set in until evaporation and is equal to
@@
From the inspection of the left panel of Fig. 1, it is clear that the memory burden significantly extends the lifespan of PBHs, with the effect becoming more pronounced as $k$ increases for a fixed $q=0.5$. This extension of the PBH lifespan results in a prolonged period of early matter domination provided by the PBHs, which leads to substantial entropy injection into the thermal bath. The right panel, which shows the dynamical evolution of energy densities, clearly illustrates the timeline of PBH evolution in the early universe. Before PBH formation, the universe is in a state of radiation domination, with $\Omega_{\rm rad}\approx 1$. After PBHs form and as the universe expands, the differing scaling behaviors of the radiation energy density $\rho_{\rm rad}$ and the PBH energy density $\rho_{\rm PBH}$ lead to a transition to matter domination, as indicated by $\Omega_{\rm PBH}$ (dashed lines) increasing towards 1, while $\Omega_{\rm rad}$ (solid lines) decreases. However, similar to the standard scenario without the memory burden effect, as PBHs lose mass, their evaporation rate increases due to the inverse dependence on $M_{\rm PBH}$ in Eq. (1), eventually reaching a turning point at $x_{\rm mb}=T_{\rm form}/T_{\rm mb}$, where $\Omega_{\rm rad}$ begins to rise again. In contrast, due to the memory burden effect, once the PBH mass reaches a threshold value $M_{\text{mb}}=qM_{\text{\rm in}}$ and the memory burden effect sets in, $\Omega_{\rm rad}$ begins to decrease once more. Thus, the key impact of the memory burden effect is to delay the return to radiation domination, effectively prolonging the period of matter domination. [^1] In principle, for sufficiently small values of $q\,(\ll 0.5)$, another intermediate radiation domination phase could arise starting around $T_{\rm mb}$ (see Appendix A). While our primary results focus on $q=0.5$, we will show that given a uniform prior for $q\in[0.2,0.8]$, the PTA fit is not so sensitive to $q$. However, for smaller values of $q$, the GWs spectrum exhibits a kink-like feature exclusive to memory-burdened PBHs. The “second phase” of matter domination, which begins around $x_{\rm mb}$, is influenced by the residual PBH mass $M_{\rm mb}$ and is much more sensitive to $k$.
@@ III Imprint of modified PBH evaporation dynamics on BGWs from inflation
re-enter the horizon at $T_{\rm eq}$ (the standard matter-radiation equality temperature), at $T_{\rm ev}$ (final evaporation temperature of memory-burdened PBHs), at $T_{\rm dom}$ (the temperature at which the PBH start to dominate the energy density) and at $T_{R}$ (when the universe reheats after inflation), respectively. The last three temperatures correspond to three characteristic frequencies describing a double peak GWs signal spaning nHz-Hz range (see, e.g. Fig. 3). The explicit expressions for those frequencies can be derived as
@@ IV Results
Here we identify the region of parameter space that optimally fits the PTA data. We employ Bayesian statistics to analyze PTA data and their posterior best-fit values with maximum likelihood, including the IPTA-DR2 and NANOGrav-15yrs datasets, utilizing the publicly available code PTArcade [97] in the ceffyl [98] mode. To mitigate the pulsar’s intrinsic excess noise at high frequencies [60, 99], we include only the first 13 and 14 frequency components from the IPTA-DR2 and NANOGrav-15yrs datasets, respectively. In Fig. 2 ($q=0.5$) and Fig. 4 ($q\in[0.2,0.8]$), we have illustrated the resulting marginalized 2D posterior distributions: green for IPTA-DR2 and orange for NANOGrav-15yrs, using the GetDist [100]. A few remarks on PTA fits are in order. The PTA fit shows limited sensitivity to $\beta$ and the reheating temperature $T_{R}$. Consideration of their appropriate values nonetheless remains essential to produce GW spectrum at higher frequencies consistent with LIGO-O3 bound and BBN constraint on $\Delta N_{\rm eff}$, where the latter is saturated by the GW amplitude at $f_{\rm high}^{\rm peak}\sim T_{R}$. To assess the sensitivity of the memory-burdened PBH parameters to PTA data, it is sufficient to impose the constraints $t_{\rm ev}\leq t_{\rm BBN}$ and $h^{2}\Omega_{\rm GW}(f_{\rm low}^{\rm peak})<5.6\times 10^{-6}\Delta N_{\rm eff}$. With these considerations, we now provide detailed discussions for two cases: first fixing $q=0.5$ which is the standard assumption in the literature on memory burden effects in black holes, and secondly varying $q$ within $[0.2,0.8]$ to see how it impacts the results.
@@
EMD provided by the memory-burdened PBHs can provide a break in the power law for frequencies corresponding to PBH evaporation, $f\sim f_{\rm low}^{\rm peak}$. More precisely, at $f_{\rm low}^{\rm peak}$, the BGWs undergo a spectral turnover from a frequency scaling of $f^{n_{T}}$ (for $f<f_{\rm low}^{\rm peak}$) to $f^{n_{T}-2}$ (for $f>f_{\rm low}^{\rm peak}$). At the same time memory-burdened PBHs with sufficiently large $M_{\rm PBH}$ or large $k$ exhibit significantly longer lifetimes, shifting $f_{\rm low}^{\rm peak}$ towards much lower frequencies within the PTA band. One might expect an early enough break in the spectrum could then allow for a good fit of the NANOGrav data for $n_{T}>3$, with the GW spectrum being less steep for $f>f_{\rm low}^{\rm peak}$. However we find that even if $T_{\rm ev}\simeq T_{\rm BBN}$, which corresponds to the minimum allowed value of $f_{\rm low}^{\rm peak}$, this frequency is too high to allow the GW spectrum to fit the NANOGrav data for $f>f_{\rm low}^{\rm peak}$. Instead, the GW spectrum remains too steep and the amplitude overshoots the PTA signal, so the $2\sigma$ upper limit $n_{T}\simeq 3$ does not change significantly compared to the standard power-law fit.
@@
Note that standard PBHs without memory burden effect [101] are obtained in the limit $k\rightarrow 0$. For comparison, we show this fit in Appendix B, where the fit favors slightly higher values of $n_{T}$ and lower $r$ and very large PBH masses are needed to increase the lifetime sufficiently. The memory burden effect with non-zero $k$ also increases the lifetime, allowing longer lifetimes for smaller PBH masses and extending the range of PBH masses that can fit the PTA data down to around $1$ g  [^2] This lower bound $M_{\rm PBH}\gtrsim 1$g arises from the maximum possible reheating temperature $T_{R}^{\rm max}\sim 5\times 10^{15}$ GeV.. Thus reducing the PBH masses can be compensated by increasing $k$ leading to the rough $k\sim 1/M_{\text{PBH}}$ correlation seen in the bottom row, third panel of Fig. 2.
While it seems from Fig. 2 that the PTA data fits allow for a large range of PBH masses, it does not account for the BBN constraint on $\Delta N_{\rm eff}$ and also LIGO-O3 bounds on higher frequencies. As mentioned previously, to describe a viable GW spectrum across a wide frequency range, the parameters $\beta$ and $T_{R}$ are crucial. In the top panels (left: NANOGrav-15yrs, right: IPTA-DR2) of Fig. 3, we show the allowed regions on the $M_{\rm PBH}$-$\beta$ plane for the minimally allowed value of $T_{R}\simeq T_{\rm form}$. The white region is consistent with PTA data as well as BBN and LIGO-O3 bounds. These plots show that a viable wide-spanning GW signal can only be obtained for a constrained range of PBH mass. It is, however, remarkable that even PBHs as light as $10^{2}$ g can leave distinct imprints on a strong wide-spanning GW signal owing to the memory burden effect. In the bottom panels, we show the GW spectrum for the best-fit values. We have chosen $\beta=10^{-0.5}$ and $\beta=10^{-3}$ for NANOGrav-15yrs and IPTA-DR2, respectively.
@@ IV.2 The case with varying $q$
We now turn our attention towards another interesting possibility with variable $q$. For phenomenological purposes we vary it from 0.2 to 0.8, thus having an additional handle over the interplay of PBH domination and PTA data. The uniform priors employed in the MCMC analysis, with the additional parameter $q$, are reported in Tab. 2. Some important points are clear from the posterior distributions shown in Fig. 4. First, there exists a very mild positive correlation between $q$ and the inflationary parameter $r$, and consequently a negative correlation between $q$ and $n_{T}$. The reason is the same as described in the previous case: as $q$ increases, the residual energy density after semiclassical evaporation rises, leading to a prolonged period of PBH domination. While $q$ does influence inflationary observables, its impact is less pronounced than that of the parameters $k$ or $M_{\rm PBH}$. Consequently, we observe a modest effect on $n_{T}$ and $r$ when analysing the PTA data. Furthermore, increasing $q$ leads to a longer PBH lifetime, causing heavier PBHs to be excluded by the $t_{\rm ev}\leq t_{\rm BBN}$ bound. Conversely, while a scenario with a lower $q$ with our chosen prior range for $k$ allows for significantly higher values, increasing $q$ can lead to conflicts with a larger $k$. This arises because both $k$ and $q$ influence the PBH lifetime in a similar manner, resulting in constraints imposed by the $t_{\rm ev}\leq t_{\rm BBN}$ at their extreme values.
The description of Fig. 5 remains the same as the one of Fig. 3, apart from the kink in the middle of the GW spectrum. This happens owing to an intermediate radiation domination for $q<0.5$. This characteristic kink feature which is unique to memory burdened PBHs, described analytically as
@@
should be present when $q<0.5$, provided that the semiclassical computation of the PBH lifetime remains valid in this regime.
Before concluding, several remarks are warranted. Our analysis relies on a power-law parameterisation of the tensor power spectrum as defined in Eq. (13). Alternative parameterisations [102, 103] may result in different phenomenological outcomes. Additionally, since we did not assume any specific inflationary model, we have neglected the running of the spectral index [104]. Our results also show sensitivity to the reheating temperature $T_{R}$. In Figs. 3 and 5, we present results for $T_{R}=T_{\rm form}$. However, we observe that as $T_{R}$ increases, the allowed parameter regions shrink (see Fig. 6). Although the predicted GW spectra have a high signal-to-noise ratio (SNR), a thorough spectral shape reconstruction, such as a Fisher analysis [105, 106], is required to assess the detectability of subtle spectral features, like $f_{\rm mb}^{\rm kink}$, with confidence. Lastly, it is important to note that PBHs may produce additional GW sources [71, 72, 73, 35, 107, 108]; however, a comprehensive study of these contributions lies beyond the scope of this article.

### v2

@@ I Introduction
The theoretical and phenomenological framework outlined above relies on a semiclassical treatment of black-hole evaporation, which has been recently challenged [40, 41, 42]. Hawking’s semiclassical calculations for 4-dimensional black holes are expected to be invalid when the black hole has lost approximately half of its initial mass. This is primarily due to the neglection of a back-reaction, wherein the quantum information retained in the system’s memory (or an excess of entropy) exerts a reciprocal effect on the black hole. This effect, also known as the “memory burden”, leads to a significant extension of the black hole’s lifetime due to the accumulation of this quantum information or memory within the black hole. Such a pivotal deviation of the PBH evaporation from the semiclassical approximation yields profound phenomenological consequences which have been investigated in the context of dark matter [43, 44, 45, 46, 47, 48, 49, 50] and primordial GWs sourced by PBHs [51, 52, 53, 54, 55, 56, 49, 57].
@@ II Impact of memory burden on PBH evaporation dynamics
where $\mathcal{A}=\mathcal{G}\,g_{*}(T_{\rm PBH})/({30720\pi})$ with $\mathcal{G}\simeq 3.8$ being the greybody factor, $g_{*}(T_{\rm PBH})\simeq 108$ is the number of relativistic degrees of freedom below the Hawking temperature $T_{\rm PBH}$, and $S(M_{\rm PBH})=4\pi M_{\rm PBH}^{2}/M_{\rm Pl}^{2}$ with $M_{\rm Pl}=1.22\times 10^{19}~{}{\rm GeV}$ is the black-hole entropy. When memory burden sets in at $t_{\rm mb}=(1-q^{3})t_{\rm ev}^{\rm sc}$, where $t_{\rm ev}^{\rm sc}=M_{\rm PBH}^{3}/(3\mathcal{A}\,M_{\rm Pl}^{4})$ is the semiclassical evaporation time, the evaporation dynamics starts to be modified allowing for a prolonged PBH lifetime before complete evaporation [^1] A precise determination of the final fate of memory-burdened PBHs requires a fully quantum gravitational treatment, which remains beyond the reach of current theoretical methods. Nevertheless, given the possibility for classical instability or continued evaporation at a suppressed rate [41], assuming full evaporation provides a well-motivated working hypothesis for the purposes of our analysis.. The total evaporation time $t_{\rm ev}$ can then be split up into,
@@
where $\tilde{t}_{\rm ev}$ is the time from when memory burden effects set in until evaporation, and is equal to
@@
From the inspection of the left panel of Fig. 1, it is clear that the memory burden significantly extends the lifespan of PBHs, with the effect becoming more pronounced as $k$ increases for a fixed $q=0.5$. This extension of the PBH lifespan results in a prolonged period of early matter domination provided by the PBHs, which leads to substantial entropy injection into the thermal bath. The right panel, which shows the dynamical evolution of energy densities, clearly illustrates the timeline of PBH evolution in the early universe. Before PBH formation, the universe was in a state of radiation domination, with $\Omega_{\rm rad}\approx 1$. After PBHs form and as the universe expands, the differing scaling behaviors of the radiation energy density $\rho_{\rm rad}$ and the PBH energy density $\rho_{\rm PBH}$ lead to a transition to matter domination, as indicated by $\Omega_{\rm PBH}$ (dashed lines) increasing towards 1, while $\Omega_{\rm rad}$ (solid lines) decreases. However, similar to the standard scenario without the memory burden effect, as PBHs lose mass, their evaporation rate increases due to the inverse dependence on $M_{\rm PBH}$ in Eq. (1), eventually reaching a turning point at $x_{\rm mb}=T_{\rm form}/T_{\rm mb}$, where $\Omega_{\rm rad}$ begins to rise again. In contrast, due to the memory burden effect, once the PBH mass reaches a threshold value $M_{\text{mb}}=qM_{\text{\rm in}}$ and the memory burden effect sets in, $\Omega_{\rm rad}$ begins to decrease once more. Thus, the key impact of the memory burden effect is to delay the return to radiation domination, effectively prolonging the period of matter domination. [^2] In principle, for sufficiently small values of $q\,(\ll 0.5)$, another intermediate radiation domination phase could arise starting around $T_{\rm mb}$ (see Appendix A). While our primary results focus on $q=0.5$, we will show that given a uniform prior for $q\in[0.2,0.8]$, the PTA fit is not so sensitive to $q$. However, for smaller values of $q$, the GW spectrum exhibits a kink-like feature exclusive to memory-burdened PBHs. The “second phase” of matter domination, which begins around $x_{\rm mb}$, is influenced by the residual PBH mass $M_{\rm mb}$ and is much more sensitive to $k$.
@@ III Imprint of modified PBH evaporation dynamics on BGWs from inflation
re-enter the horizon at $T_{\rm eq}$ (the standard matter-radiation equality temperature), at $T_{\rm ev}$ (final evaporation temperature of memory-burdened PBHs), at $T_{\rm dom}$ (the temperature at which the PBH start to dominate the energy density) and at $T_{R}$ (when the universe reheats after inflation), respectively. The last three temperatures correspond to three characteristic frequencies describing a double peak GW signal spanning the nHz-Hz range (see, e.g. Fig. 3). The explicit expressions for those frequencies can be derived as
@@ IV Results
Here we identify the region of parameter space that optimally fits the PTA data. We employ Bayesian statistics to analyze PTA data and their posterior best-fit values with maximum likelihood, including the IPTA-DR2 and NANOGrav-15yrs datasets, utilizing the publicly available code PTArcade [97] in the ceffyl [98] mode. To mitigate the pulsar’s intrinsic excess noise at high frequencies [60, 99], we include only the first 13 and 14 frequency components from the IPTA-DR2 and NANOGrav-15yrs datasets, respectively. In Fig. 2 ($q=0.5$) and Fig. 4 ($q\in[0.2,0.8]$), we have illustrated the resulting marginalized 2D posterior distributions: green for IPTA-DR2 and orange for NANOGrav-15yrs, using the GetDist [100]. A few remarks on PTA fits are in order. The PTA fit shows limited sensitivity to $\beta$ and the reheating temperature $T_{R}$. Consideration of their appropriate values nonetheless remains essential to produce GW spectrum at higher frequencies consistent with LIGO-O3 bound and BBN constraint on $\Delta N_{\rm eff}$, where the latter is saturated by the GW amplitude at $f_{\rm high}^{\rm peak}\sim T_{R}$. To assess the sensitivity of the memory-burdened PBH parameters to PTA data, it is sufficient to impose the constraints $t_{\rm ev}\leq t_{\rm BBN}$ and $h^{2}\Omega_{\rm GW}(f_{\rm low}^{\rm peak})<5.6\times 10^{-6}\Delta N_{\rm eff}$. With these considerations, we now provide detailed discussions for two cases: first, fixing $q=0.5$, which is the standard assumption in the literature on memory burden effects in black holes, and secondly, varying $q$ within $[0.2,0.8]$ to see how it impacts the results.
@@
EMD provided by the memory-burdened PBHs can provide a break in the power law for frequencies corresponding to PBH evaporation, $f\sim f_{\rm low}^{\rm peak}$. More precisely, at $f_{\rm low}^{\rm peak}$, the BGWs undergo a spectral turnover from a frequency scaling of $f^{n_{T}}$ (for $f<f_{\rm low}^{\rm peak}$) to $f^{n_{T}-2}$ (for $f>f_{\rm low}^{\rm peak}$). At the same time, memory-burdened PBHs with sufficiently large $M_{\rm PBH}$ or large $k$ exhibit significantly longer lifetimes, shifting $f_{\rm low}^{\rm peak}$ towards much lower frequencies within the PTA band. One might expect an early enough break in the spectrum could then allow for a good fit of the NANOGrav data for $n_{T}>3$, with the GW spectrum being less steep for $f>f_{\rm low}^{\rm peak}$. However we find that even if $T_{\rm ev}\simeq T_{\rm BBN}$, which corresponds to the minimum allowed value of $f_{\rm low}^{\rm peak}$, this frequency is too high to allow the GW spectrum to fit the NANOGrav data for $f>f_{\rm low}^{\rm peak}$. Instead, the GW spectrum remains too steep and the amplitude overshoots the PTA signal, so the $2\sigma$ upper limit $n_{T}\simeq 3$ does not change significantly compared to the standard power-law fit.
@@
Note that standard PBHs without memory burden effect [101] are obtained in the limit $k\rightarrow 0$. For comparison, we show this fit in Appendix B, where the fit favors slightly higher values of $n_{T}$ and lower $r$ and very large PBH masses are needed to increase the lifetime sufficiently. The memory burden effect with non-zero $k$ also increases the lifetime, allowing longer lifetimes for smaller PBH masses and extending the range of PBH masses that can fit the PTA data down to around $1$ g  [^3] This lower bound $M_{\rm PBH}\gtrsim 1$g arises from the maximum possible reheating temperature $T_{R}^{\rm max}\sim 5\times 10^{15}$ GeV.. Thus, reducing the PBH masses can be compensated by increasing $k$, leading to the rough $k\sim 1/M_{\text{PBH}}$ correlation seen in the bottom row, third panel of Fig. 2.
While it seems from Fig. 2 that the PTA data fits allow for a large range of PBH masses, it does not account for the BBN constraint on $\Delta N_{\rm eff}$ and also LIGO-O3 bounds on higher frequencies. As mentioned previously, to describe a viable GW spectrum across a wide frequency range, the parameters $\beta$ and $T_{R}$ are crucial. In the top panels (left: NANOGrav-15yrs, right: IPTA-DR2) of Fig. 3, we show the allowed regions on the $M_{\rm PBH}$-$\beta$ plane for the minimally allowed value of $T_{R}\simeq T_{\rm form}$. The white region is consistent with PTA data as well as BBN and LIGO-O3 bounds. These plots show that a viable, wide-spanning GW signal can only be obtained for a constrained range of PBH mass. It is, however, remarkable that even PBHs as light as $10^{2}$ g can leave distinct imprints on a strong, wide-spanning GW signal owing to the memory burden effect. In the bottom panels, we show the GW spectrum for the best-fit values. We have chosen $\beta=10^{-0.5}$ and $\beta=10^{-3}$ for NANOGrav-15yrs and IPTA-DR2, respectively.
@@ IV.2 The case with varying $q$
We now turn our attention towards another interesting possibility with variable $q$. For phenomenological purposes, we vary it from 0.2 to 0.8, thus having an additional handle over the interplay of PBH domination and PTA data. The uniform priors employed in the MCMC analysis, with the additional parameter $q$, are reported in Tab. 2. Some important points are clear from the posterior distributions shown in Fig. 4. First, there exists a very mild positive correlation between $q$ and the inflationary parameter $r$, and consequently a negative correlation between $q$ and $n_{T}$. The reason is the same as described in the previous case: as $q$ increases, the residual energy density after semiclassical evaporation rises, leading to a prolonged period of PBH domination. While $q$ does influence inflationary observables, its impact is less pronounced than that of the parameters $k$ or $M_{\rm PBH}$. Consequently, we observe a modest effect on $n_{T}$ and $r$ when analysing the PTA data. Furthermore, increasing $q$ leads to a longer PBH lifetime, causing heavier PBHs to be excluded by the $t_{\rm ev}\leq t_{\rm BBN}$ bound. Conversely, while a scenario with a lower $q$ with our chosen prior range for $k$ allows for significantly higher values, increasing $q$ can lead to conflicts with a larger $k$. This arises because both $k$ and $q$ influence the PBH lifetime in a similar manner, resulting in constraints imposed by the $t_{\rm ev}\leq t_{\rm BBN}$ at their extreme values.
The description of Fig. 5 remains the same as that of Fig. 3, apart from the kink in the middle of the GW spectrum. This happens owing to an intermediate radiation domination for $q<0.5$. This characteristic kink feature, which is unique to memory-burdened PBHs, is described analytically as
@@
and should be present when $q<0.5$, provided that the semiclassical computation of the PBH lifetime remains valid in this regime.
Before concluding, several remarks are warranted. Our analysis relies on a power-law parameterisation of the tensor power spectrum as defined in Eq. (13). Alternative parameterisations [102, 103] may result in different phenomenological outcomes. Additionally, since we did not assume any specific inflationary model, and as shown in ref. [104], the minimal Constant Power Law (CPL) model remains sufficient to describe the current NANOGrav signal. While future datasets may allow for a meaningful measurement of nonzero spectral running, at present, its role at PTA frequencies is marginal and does not affect our numerical results regarding PTA constraints on the PBH parameter space. Therefore, we have neglected the running of the spectral index [105]. For completeness, we provide a discussion of its effects in Appendix C, where we explicitly show how it modifies the SGWB spectrum over a wide frequency range, but this has no impact on our main findings. Our results also show sensitivity to the reheating temperature $T_{R}$. In Figs. 3 and 5, we present results for $T_{R}=T_{\rm form}$. However, we observe that as $T_{R}$ increases, the allowed parameter regions shrink (see Fig. 6). Although the predicted GW spectra have a high signal-to-noise ratio (SNR), a thorough spectral shape reconstruction, such as a Fisher analysis [106, 107], is required to assess the detectability of subtle spectral features, like $f_{\rm mb}^{\rm kink}$, with confidence. Lastly, it is important to note that PBHs may produce additional GW sources [71, 72, 73, 35, 108, 109]; however, a comprehensive study of these contributions lies beyond the scope of this article.